# Dissecting the NVIDIA Hopper Architecture through Microbenchmarking and Multiple Level Analysis

### v1

@@ 4.2 Throughput
For the global memory test, we allocate significantly more memory space than the capacity of L2 to bypass the hardware-level cache prefetch mechanism, as detailed in [28]. In order to reduce the number of memory access instructions, we set up each thread to use vectorized memory access to read float4. Each thread reads 5 times and writes 1 time. The number of blocks we used is four times the number of SMs.
For the first three units, data is accessed as both single-precision floating-point values (float) and four-element vectors (float4). While the float type is more common, float4 allows for data transfer with fewer instructions. For global memory accesses, we exclusively employ vectorized (float4) reads and writes to minimize instruction overhead and maximize memory bandwidth utilization [41]. We record the cycles or time consumed and the amount of data accessed to calculate the throughput of different units.
Table V shows the memory access throughput at different memory levels. In the cache throughput test, we use different data types for memory access. We observe that using vectorized memory access (FP32.v4, equivalent to CUDA’s float4) can always achieve better performance. Notably, on the consumer-grade RTX 4090, using FP32 to access L1 cache and shared memory yields only half the throughput achieved with FP32.v4. This performance discrepancy is attributed to memory input/output throttling on the RTX 4090. Utilizing wider data types or increasing the number of blocks can mitigate this bottleneck.
The maximum throughput of L1 cache and shared memory of the three devices are similar because they share same hardware. However, in terms of L2 cache throughput, the H800 is 2.6 times and 2.2 times higher than the RTX 4090 and A100, respectively. [^1] Note that for the L1 cache, the amount of data they transfer per clock is almost the same. However, since the order of clock frequency from high to low is RTX4090, H800, and A100, the order of throughput per unit time from high to low is also RTX4090, H800, and A100. The same calculation method also applies to L2 cache. In the memory throughput test, our results reach 92%, 90%, and 91% of the theoretical performance on RTX4090, A100, and H800 respectively. In the comparison of L2 and Global, the L2 cache throughput of RTX4090, A100, and H800 is 4.67, 2.01, and 4.23 times the global memory throughput, respectively.
TABLE V: Throughput at different memory levels
@@
L1 Cache (byte/clk/SM) FP32 FP32.v4 FP32 FP32.v4 FP32 FP32.v4
63.7 121.2 99.5 106.8 125.8 124.1
L2 Cache (byte/clk) FP32 FP32.v4 FP32 FP32.v4 FP32 FP32.v4
1622.2 1708.0 1853.7 2007.9 4472.3 3942.4
Shared Memory (byte/clk/SM) FP32 FP32.v4 FP32 FP32.v4 FP32 FP32.v4
63.7 126.5 127.9 126.2 127.4 127.9
Global Memory (GB/s) 929.8 1407.2 1861.5
@@
5 Tensor Memory Accelerator
The asynchronous copies introduced by Nvidia Ampere GPU architecture provide more flexibility for programming. Hopper provides a more sophisticated asynchronous copy engine: the Tensor Memory Accelerator (TMA). As shown in Fig. 1, TMA can be used in two scenarios: (1) global memory and shared memory transfer in both directions. (2) shared memory regions of different SMs within a cluster transfer (distributed shared memory). Ampere’s asynchronous copies require all threads to calculate the address of their own memory accesses. However, Hopper’s TMA takes care of everything. We can assign one thread to create a copy descriptor before launching the TMA. After that, TMA can automatically handle addressing and data transfer.
@@ 5.2 Throughput
Figure 3: Global memory throughput of different combinations of block numbers and TMA load sizes
@@
We compare the performance of different shared memory box shapes with a TMA load size of 16KB and floating-point data type. The configurations tested are 16×16×16, 64×64×1, 256×16×1, 16×256×1, and 4×4×256. The results, shown in Fig. 4, reveal a clear trend: with the same load size, larger x-axis dimensions yield higher throughput without negative effects from increasing thread blocks. However, increasing the y-axis and z-axis dimensions significantly reduces throughput, which may decrease further with more thread blocks. Therefore, when using TMA to handle tensors with dimensions greater than 3D, it’s important to carefully select appropriate parameters to achieve optimal performance.
Figure 4: 16KB per TMA load with different shapes
6 Tensor Core
6.1 Tensor Core’s Evolution
@@
Table VI illustrates the progression of TCs, encompassing enhancements in precision, operand shapes, programming modes, and execution modes. In the initial Volta Architecture, first-generation TCs exclusively supported FP16 as the input data type. Subsequent architectures, including Ampere, Ada, and Hopper, introduced support for a broader range of data types such as BF16, TF32, FP64, INT8, INT4, Binary, and more. The programming of TCs has also seen continuous improvement. Ampere and Ada Lovelace GPUs provide users with the flexibility to utilize either the legacy C-level wmma APIs or PTX-level mma instructions. Notably, the wmma APIs had limitations in fully harnessing TCs’ capabilities, whereas mma instructions could leverage advanced sparse matrix multiplication capabilities introduced since Ampere. In the case of Hopper GPUs, new warp-group-level wgmma instructions were introduced. Both wmma and mma APIs remain supported in Hopper, but we find that the complete potential of Hopper TCs can only be realized through wgmma instructions.
@@
Fig. 5 provides examples of both mma and wgmma instructions, demonstrating mixed-precision capabilities. An mma instruction computes $\mathbf{D}_{m\times n}=\mathbf{A}_{m\times k}\times\mathbf{B}_{k\times n}+\mathbf{C}_{m\times n}$ and is executed synchronously by one CUDA warp (i.e., 32 threads). In contrast, wgmma for Hopper computes $\mathbf{D}_{m\times n}=\mathbf{A}_{m\times k}\times\mathbf{B}_{k\times n}+\mathbf{D}_{m\times n}$ and is executed asynchronously by one CUDA warp group (i.e., four CUDA warps). Here, $\mathbf{A}_{m\times k}$ denotes a matrix $\mathbf{A}$ with $m$ rows and $k$ columns, and similarly for the other matrices. The matrix shapes for mma instructions can be $m16n8k16$ or $m16n8k8$, while wgmma supports $m64nNk16$, where $N$ can be 16, 32, 64, 128, 256, and so on [42]. Notably, wgmma has the advantage of directly loading matrices $A$ and $B$ from shared memory, unlike mma, which requires storing all matrices in the register file before execution. We use the term “SS” to denote the wgmma instruction that loads both $A$ and $B$ from shared memory, while “RS” is used for the instruction that loads $A$ from the register file. Additionally, wgmma offers support for certain useful arguments not required for mma. Further details are provided in [42].
TABLE VI: Properties of the latest generations of Tensor Cores
@@
6.2 Instruction-level
We conduct micro-benchmarking of TCs at the PTX level, as it strikes a suitable balance between granularity and complexity. Additionally, we disassemble PTX instructions to SASS codes to achieve a deeper understanding of the operations.
Figure 5: The mma and wgmma instructions that perform $D=A\times B+C$ and $D=A\times B\{+D\}$, respectively.
[…]
SASS analysis. We perform the disassembly of mma and wgmma instructions specifically for Hopper Tensor Cores, and the results are presented in Table VII. The mma instructions undergo compilation into SASS instructions, with the naming convention following the established patterns: HMMA (for floating-point types), IMMA (for integer types), and BMMA (for binary types). Notably, there are two specialized types within mma: INT4 and FP8.
TABLE VII: SASS Instructions for Different Hopper Tensor Core PTX Instructions
A/B C/D mma wgmma
FP16 FP16 HMMA.16816.F16 HGMMA.64x256x16.F16
FP16 FP32 HMMA.16816.F32 HGMMA.64x256x16.F32
TF32 FP32 HMMA.1688.F32.TF32 HGMMA.64x256x8.F32.TF32
FP8 FP16 $\times$
QGMMA.64x256x32.F16.E5M2.E5M2
QGMMA.64x256x32.F16.E4M3.E4M3
FP8 FP32 HMMA.16816.F32
QGMMA.64x256x32.F32.E4M3.E4M3
QGMMA.64x256x32.F32.E5M2.E5M2
INT8 INT32 IMMA.16832.S8.S8 IGMMA.64x256x32.S8.S8
INT4 INT32
IMAD.MOV.U32
$\times$
Binary INT32 BMMA.168256.AND.POPC BGMMA.64x256x256.AND.POPC
mma results. Table VIII provides an overview of the latency and throughput measurements for mma instructions across A100, RTX4090, and H800 Tensor Cores GPUs. Note that the sparse shapes in the table represent compressed shapes. In other words, the $k$ of the actual instruction modifier is twice that in the table.
TABLE VIII: Different dense and sparse mma instructions on A100, RTX4090 and H800 Tensor Cores. Latency (LAT) is measured in clock cycles. Throughput is measured in TFLOPS or TOPS/s. Peak performance (A100): FP16 (312 TFLOPS); TF32 (156 TFLOPS); INT8 (624 TOPS). Peak performance (RTX4090): FP16 (330.3 TFLOPS); TF32 (82.6 TFLOPS); INT8 (660.6 TOPS). Peak performance (H800): FP16 (756.5 TFLOPS); TF32 (378 TFLOPS); INT8 (1513 TOPS).
@@ 6.2 Instruction-level
wgmma results. As a set of warp-group-level Tensor Core instructions designed specifically for Hopper GPUs, wgmma instructions are the pioneering instructions to be executed asynchronously. Table IX and Table X show the measured latency and throughput of dense and sparse wgmma instructions, respectively. When initializing matrices with zeros, we achieve throughputs exceeding 95% of the theoretical peak performance. We observe a decrease in Tensor Core performance when initializing matrices with random values, especially pronounced when utilizing FP16 as the computation type and FP32 for accumulation. This phenomenon is primarily attributed to the power consumption nearing the 350 W power limit of the H800-PCIe, subsequently causing a reduction in frequency. Users working with Tensor Cores on the H800-PCIe GPU should take into full consideration the power constraints when performing computations.
In the context of sparse wgmma instructions, the latencies for “RS” and “SS” modes are 128.0 and 144.0, respectively. Additionally, we observe that the achieved throughputs in “SS” modes are lower compared to “RS” modes, which is notably different from dense wgmma. We find that in sparse wgmma, the “SS” mode retrieves data from the shared memory of size $m\times k$ and performs a 2:4 sparse pruning based on metadata during the execution of the sparse wgmma instruction. In contrast, the “RS” mode directly accesses data from the pruned register file of size $m\times k/2$. The high shared memory access demand (twice as much) may lead to latency that cannot be effectively concealed by Tensor Core computation, resulting in sparse wgmma instructions in “SS” modes failing to achieve the expected peak performance.
TABLE IX: Variations in Dense wgmma Instructions for H800 Tensor Cores. Latency (LAT) is quantified in clock cycles, while throughput is expressed in TFLOPS or TOPS. The peak throughputs for FP16, TF32, FP8, and INT8 are 756.5, 373, 1513, and 1513, correspondingly. “Zero” or “Rand” signifies that all matrices are initialized with either zero or randomly generated values. “SS” implies that both matrix A and B are stored in shared memory, while “RS” signifies that matrix A is stored in the register file, whereas B is stored in shared memory. “Rand” has the same latency as “Zero”.
@@
TABLE X: Different sparse wgmma Instructions on H800 Tensor Cores. The definitions of LAT and throughput can be found in the caption of Table IX. “Rand” has the same latency as “Zero”.
A/B C/D
Sparse
Instruction
LAT/Throughput
(SS,Zero)
LAT/Throughput
@@
Throughput
(SS,Rand)
Throughput
@@
FP16 FP16 sp.m64n256k32 144.0/1308.0 128.0/1472.0 1257.8 1362.3
FP16 FP32 sp.m64n256k32 144.0/1312.3 128.0/1476.2 1194.3 1277.5
TF32 FP32 sp.m64n256k16 144.0/656.8 128.0/735.4 644.9 721.7
FP8 FP16 sp.m64n256k64 144.0/2619.9 128.0/2945.0 2588.6 2782.4
FP8 FP32 sp.m64n256k64 144.0/2622.8 128.0/2931.0 2588.7 2722.3
INT8 INT32 sp.m64n256k64 144.0/2612.4 128.0/2933.0 2593.9 2898.3
wgmma results with different $N$ values. We conduct tests using the example of wgmma.m64nNk16.f32.f16.f16, varying the value of $N$, and the results are presented in Table XI. When $N$ is greater than or equal to 64, all wgmma instructions can achieve throughputs that closely approach peak performance. However, when $N$ is less than 64, the achieved throughput decreases, and the “SS” mode of instructions exhibits higher latency than the “RS” mode, while the achieved throughputs are lower than those of the “RS” mode. As $N$ decreases, the computational density of wgmma instructions gradually diminishes, making it challenging to conceal the latency associated with shared memory access, leading to the aforementioned phenomena. Therefore, when utilizing wgmma instructions, it is advisable to opt for larger values of $N$ ($>=64$) whenever possible to attain superior performance.
TABLE XI: Different wgmma instructions with different $N$ values on H800 tensor cores. The definitions of LAT and throughput can be found in the caption of Table IX. “Rand” has the same latency as “Zero”.
N Dense Sparse
LAT/Throughput
(SS,Zero)
LAT/Throughput
@@
Throughput
(SS,Rand)
Throughput
(RS,Rand)
LAT/Throughput
(SS,Zero)
LAT/Throughput
(RS,Zero)
Throughput
(SS,Rand)
Throughput
@@
Energy efficiency. In our benchmarks, we define energy efficiency as performance per watt, measured in TFLOPS(TOPS)/watt. Both mma and wgmma are considered in this work. Our methodology involves repeatedly executing either mma or wgmma instructions across all SMs, with a minimum of two billion repetitions. We utilize the tools provided in [16] to monitor both frequency and performance. The energy efficiency of each instruction is determined by calculating the ratio of performance to power consumption once steady-state operation is achieved during these repetitive executions.
@@
For mma instructions, we test the largest operational shape in Table VIII. The energy efficiency of mma instructions is shown in Table XII. In terms of dense instructions, the average energy efficiency of H800 is 1.60 times and 1.69 times that of A100 and RTX4090 respectively. In terms of sparse instructions, the average energy efficiency of H800 is 1.33 times and 1.39 times that of A100 and RTX4090 respectively. We find that the H800 has significantly higher energy efficiency.
@@
TABLE XII: Power consumption and energy efficiency of maximum shape under mma instructions. T, D, and S represent Type, Dense, and Sparse respectively. P stands for energy, measured in Watts. E stands for efficiency, measured in TFLOPS(TOPS)/watt.
A/B C/D T A100 H800 RTX4090
P E P E P E
FP16 FP16 D 173.4 1.79 188.6 2.62 189.1 1.89
S 198.8 3.13 187.2 3.86 214.0 3.33
FP16 FP32 D 188.5 1.61 196.7 2.49 154.1 1.16
S 216.1 2.79 194.9 3.70 165.9 2.15
TF32 FP32 D 214.7 0.71 254.9 0.97 174.3 0.51
S 235.7 1.28 232.5 1.56 187.9 0.95
INT8 INT32 D 178.4 3.41 165.3 5.92 201.4 3.53
S 193.9 6.24 163.3 8.79 219.8 6.47
@@
For wgmma instructions, we focus on both the frequency and efficiency when these instructions are executed continuously. The instruction formats are consistent with those described in Tables IX and X. The results are presented in Table XIII. It is observed that, during continuous execution of wgmma instructions, the core frequency drops below 1620 MHz specified in the whitepaper, as it reaches the 350-watt power limit. Furthermore, the initialization of inputs significantly impacts both power consumption and frequency. When inputs are set to zero, power consumption remains below 200 watts. However, with random inputs, the power consumption rapidly hits the 350-watt threshold. Consequently, the performance results in Table XIII are inferior to those in Tables IX and X. Notably, sparse instructions cause a substantial reduction in frequency, explaining why their performance is not twice that of dense instructions. Regarding energy efficiency, wgmma dense and sparse instructions achieve average 0.67 times and 0.78 times the efficiency of mma dense and sparse instructions, respectively. Compared to different architectures, the energy efficiency of Hopper wgmma is on average 1.05 times and 1.10 times that of Ampere mma and Ada mma instructions, respectively.

### v2

@@ 4.2. Throughput
For the global memory test, we allocate significantly more memory space than the capacity of L2 to bypass the hardware-level cache prefetch mechanism, as detailed in (Mei and Chu, 2017). In order to reduce the number of memory access instructions, we set up each thread to use vectorized memory access to read float4. Each thread reads 5 times and writes 1 time. The number of blocks we used is four times the number of SMs. The number of blocks has little effect on L2 performance testing, requiring only enough instructions to saturate the hardware units.
For the first three units, data is accessed as both single-precision floating-point values (float) and four-element vectors (float4). While the float type is more common, float4 allows for data transfer with fewer instructions. For global memory accesses, we exclusively employ vectorized (float4) reads and writes to minimize instruction overhead and maximize memory bandwidth utilization (Luitjens, 2013). We record the cycles or time consumed and the amount of data accessed to calculate the throughput of different units.
Table 5 shows the memory access throughput at different memory levels. In the cache throughput test, we use different data types for memory access. We observe that using vectorized memory access (FP32.v4, equivalent to CUDA’s float4) can always achieve better performance. Notably, on the consumer-grade RTX 4090, using FP32 to access L1 cache and shared memory yields only half the throughput achieved with FP32.v4. This performance discrepancy is attributed to memory input/output throttling on the RTX 4090. Utilizing wider data types or increasing the number of blocks can mitigate this bottleneck.
The maximum throughput of L1 cache and shared memory of the three devices are similar because they share the same hardware design. However, in terms of L2 cache throughput, the H800 is 2.6 times and 2.2 times higher than the RTX 4090 and A100, respectively.Note that for the L1 cache, the amount of data they transfer per clock is almost the same. However, since the order of clock frequency from high to low is RTX4090, H800, and A100, the order of throughput per unit time from high to low is also RTX4090, H800, and A100. The same calculation method also applies to L2 cache. In the memory throughput test, our results reach 92%, 90%, and 91% of the theoretical performance on RTX4090, A100, and H800 respectively. In the comparison of L2 and Global, the L2 cache throughput of RTX4090, A100, and H800 is 4.67, 2.01, and 4.23 times the global memory throughput, respectively.
Table 5. Throughput at different memory levels. The unit for L1 cache and shared memory throughput is (byte/clk/SM). The unit for L2 cache throughput is (byte/clk). The unit for global memory throughput is GB/s.
@@
FP32 FP32.v4 FP32 FP32.v4 FP32 FP32.v4
L1 Cache 63.7 121.2 99.5 106.8 125.8 124.1
Shared Memory 63.7 126.5 127.9 126.2 127.4 127.9
L2 Cache 1622.2 1708.0 1853.7 2007.9 4472.3 3942.4
Global Memory 929.8 1407.2 1861.5
@@
Insight 1. Hopper’s L1 cache and shared memory performance remain consistent with the previous generation. 2. The L2 cache architecture in both Ampere and Hopper features a dual-partition design, where partitions communicate through high-bandwidth interconnection channels. GPCs are similarly organized into two partitions that align with the L2 cache structure, exhibiting optimal performance when accessing proximate L2 cache partitions.
5. Tensor Memory Accelerator
The asynchronous copies introduced by NVIDIA Ampere GPU architecture provide more flexibility for programming. Hopper provides a more sophisticated asynchronous copy engine: the Tensor Memory Accelerator (TMA). As shown in Fig. 1, TMA can be used in two scenarios: (1) transfers between global memory and shared memory in both directions; (2) transfers between shared memory regions of different SMs within a cluster (distributed shared memory). Ampere’s asynchronous copies require all threads to calculate the address of their own memory accesses. However, Hopper’s TMA takes care of everything. We can assign one thread to create a copy descriptor before launching the TMA. After that, TMA can automatically handle addressing and data transfer.
@@ 5.2. Throughput
Figure 3. Global memory throughput of different combinations of block numbers and TMA load sizes
@@
We compare the performance of different shared memory box shapes with a TMA load size of 16KB and floating-point data type. The configurations tested are 16×16×16, 32x128x1, 64×64×1, 256×16×1, 16×256×1, and 4×4×256. The results, shown in Fig. 4, reveal a clear trend: with the same load size, larger x-axis dimensions yield higher throughput without negative effects from increasing thread blocks. However, increasing the y-axis and z-axis dimensions significantly reduces throughput, which may decrease further with more thread blocks. Therefore, when using TMA to handle tensors with dimensions greater than 3D, it’s important to carefully select appropriate parameters to achieve optimal performance.
Figure 4. 16KB per TMA load with different shapes
5.3. Application-level
For the application-level evaluation of TMA, we select one of the most impactful and widely used applications: General Matrix Multiply (GEMM). GEMM is defined as $\mathbf{D}_{m\times n}=\alpha\mathbf{A}_{m\times k}\times\mathbf{B}_{k\times n}+\beta\mathbf{C}_{m\times n}$. GEMM is extensively applied in fields such as artificial intelligence and scientific computing. GPUs, with their high degree of parallelism, have become the preferred hardware for GEMM computations. To achieve high efficiency on GPUs, GEMM implementations employ tiling strategies, which divide the computational workload into smaller submatrices (tiles) that fit into the memory hierarchy of the GPU. Each tile is assigned to a Cooperative Thread Array (CTA), which consists of threads that collaboratively load data from global memory, perform computations in shared memory, and write results back to global memory. This tiling approach minimizes memory latency and maximizes data reuse, ensuring efficient utilization of GPU resources.
Our implementation is based on the official example provided by CUTLASShttps://github.com/NVIDIA/cutlass/tree/main/examples/cute/tutorial/hopper. To minimize the influence of other factors (e.g., Hopper’s more powerful Tensor Cores and its wgmma instructions), this subsection focuses exclusively on comparisons conducted on the Hopper architecture. In the GEMM implementation without TMA, the programming model follows the same approach as Volta and Ampere, employing pipelining by interleaving stages and instructions. In the TMA-enabled version, explicit producer-consumer synchronization is utilized for purely asynchronous instructions, such as TMA and Tensor Core operations. To maintain high performance for larger matrices, we set the CTA tile sizes for the $M$, NN, and $K$ dimensions to 128, 256, and 64, respectively. At this point, the TMA load shapes are configured as 128×256 and 64×256 (with the $\mathbf{B}$ matrix being transposed), which are equivalent to 64×256 and 32×256 in FP32. Based on the instruction-level results, these configurations allow for effective utilization of TMA’s performance capabilities. To fully utilize the performance of Tensor Cores (as discussed in Table 9 of Section 6.2, where it is shown that the $N$ dimension needs to be sufficiently large to ensure tensor core performance), we set the $M$, NN, and $K$ dimensions for Tensor Core instructions to 64, 128, and 16, respectively. Additionally, we configure the pipeline depth to 2. All other parameters remain consistent with the CUTLASS example.
Figure 5. GEMM Performance Comparison in FP16: With and Without TMA
Fig. 5 illustrates the performance impact of enabling and disabling TMA in the GEMM application. The horizontal axis represents the dimensions of square matrices $(n\times n)$, while the vertical axis shows the performance in FP16 TFLOPS (Tera Floating-point Operations Per Second). For smaller matrices $(n<2000)$, the performance difference between “With TMA” (blue line) and “Without TMA” (orange line) is minimal, as the smaller workload does not fully leverage GPU’s capabilities. As the matrix size increases $(n\geq 2000)$, the performance with TMA shows a significant advantage, growing faster than the case without TMA. For larger matrices $(n>7000)$, TMA-enabled performance stabilizes at close to 600 TFLOPS, significantly outperforming the non-TMA case, which levels off at approximately 450 TFLOPS. This demonstrates that TMA provides substantial benefits for large-scale matrix multiplication by optimizing memory access and data movement, enabling higher computational efficiency and better utilization of hardware resources.
Insight 1. TMA, as an independent asynchronous memory access unit, enhances the programming flexibility of the Hopper architecture. Utilizing warp specialization can more fully exploit the performance potential of the Hopper architecture. 2. TMA load shapes require careful selection to achieve good performance, ensuring the x-axis dimension is sufficiently large while fully considering memory-compute overlap in the pipeline.
6. Tensor Core
6.1. Tensor Core’s Evolution
@@
Table 6 illustrates the progression of TCs, encompassing enhancements in precision, operand shapes, programming modes, and execution modes. In the initial Volta Architecture, first-generation TCs exclusively supported FP16 as the input data type. Subsequent architectures, including Ampere, Ada, and Hopper, introduced support for a broader range of data types such as BF16, TF32, FP64, INT8, INT4, Binary, and more. The programming of TCs has also seen continuous improvement. Ampere and Ada Lovelace GPUs provide users with the flexibility to utilize either the legacy C-level wmma APIs or PTX-level mma instructions. Notably, the wmma APIs had limitations in fully harnessing TCs’ capabilities, whereas mma instructions could leverage advanced sparse matrix multiplication capabilities introduced since Ampere. In the case of Hopper GPUs, new warp-group-level wgmma instructions were introduced. Both wmma and mma APIs remain supported in Hopper, but we find that the complete potential of Hopper TCs can only be realized through wgmma instructions.
@@
An mma instruction computes $\mathbf{D}_{m\times n}=\mathbf{A}_{m\times k}\times\mathbf{B}_{k\times n}+\mathbf{C}_{m\times n}$ and is executed synchronously by one CUDA warp (i.e., 32 threads). In contrast, wgmma for Hopper computes $\mathbf{D}_{m\times n}=\mathbf{A}_{m\times k}\times\mathbf{B}_{k\times n}+\mathbf{D}_{m\times n}$ and is executed asynchronously by one CUDA warp group (i.e., four CUDA warps). Here, $\mathbf{A}_{m\times k}$ denotes a matrix $\mathbf{A}$ with $m$ rows and $k$ columns, and similarly for the other matrices. The matrix shapes for mma instructions can be $m16n8k16$ or $m16n8k8$, while wgmma supports $m64nNk16$, where $N$ can be 16, 32, 64, 128, 256, and so on (Corporation, 2023). Notably, wgmma has the advantage of directly loading matrices $A$ and $B$ from shared memory, unlike mma, which requires storing all matrices in the register file before execution. We use the term “SS” to denote the wgmma instruction that loads both $A$ and $B$ from shared memory, while “RS” is used for the instruction that loads $A$ from the register file. Additionally, wgmma offers support for certain useful arguments not required for mma. Further details are provided in (Corporation, 2023). Notably, wgmma instructions are specific to the Hopper architecture, whereas mma instructions offer backward and forward compatibility across generations.
Table 6. Properties of the latest generations of Tensor Cores
@@
6.2. Instruction-level
We conduct micro-benchmarking of TCs at the PTX level, as it strikes a suitable balance between granularity and complexity. Additionally, we disassemble PTX instructions to SASS codes to achieve a deeper understanding of the operations. SASS analysis is provided in the supplementary materials, which are available online.
We focus on assessing two critical metrics: latency and throughput. Latency signifies the elapsed time, measured in clock cycles, starting from the initiation of instruction issuance to the execution pipeline and concluding when the results become accessible for subsequent usage. This measurement is specifically labeled as “completion latency.” To elaborate further, we issue a single synchronous TC instruction (i.e., mma) using one CUDA warp per SM, whereas one asynchronous TC instruction (i.e., wgmma) is issued utilizing four CUDA warps (comprising a warp group) on one SM. We execute the instruction 1024 times within a CUDA kernel. Throughput is quantified as $\mathit{Total\_{OPS}}/\mathit{Duration}$, where $\mathit{OPS}$ represents multiplication or addition operations. It’s important to emphasize that, unlike the approach described in (Sun et al., 2023), we abstain from utilizing total clock cycles to compute throughput due to potential variations in GPU frequencies during the execution of different TC instructions.
mma results. Table 7 provides an overview of the latency and throughput measurements for mma instructions across A100, RTX4090, and H800 Tensor Cores GPUs. Note that the sparse shapes in the table represent compressed shapes. In other words, the $k$ of the actual instruction modifier is twice that in the table.
Table 7. Different dense and sparse mma instructions on A100, RTX4090 and H800 Tensor Cores. Latency (LAT) is measured in clock cycles. Throughput is measured in TFLOPS or TOPS/s. Peak performance (A100): FP16 (312 TFLOPS); TF32 (156 TFLOPS); INT8 (624 TOPS). Peak performance (RTX4090): FP16 (330.3 TFLOPS); TF32 (82.6 TFLOPS); INT8 (660.6 TOPS). Peak performance (H800): FP16 (756.5 TFLOPS); TF32 (378 TFLOPS); INT8 (1513 TOPS).
@@ 6.2. Instruction-level
wgmma results. As a set of warp-group-level Tensor Core instructions designed specifically for Hopper GPUs, wgmma instructions are the pioneering instructions to be executed asynchronously. Table 8 show the measured latency and throughput of dense instructions. When initializing matrices with zeros, we achieve throughputs exceeding 95% of the theoretical peak performance. We observe a decrease in Tensor Core performance when initializing matrices with random values, especially pronounced when utilizing FP16 as the computation type and FP32 for accumulation. This phenomenon is primarily attributed to the power consumption nearing the 350 W power limit of the H800-PCIe, subsequently causing a reduction in frequency. Users working with Tensor Cores on the H800-PCIe GPU should take into full consideration the power constraints when performing computations.
Table 8. Variations in Dense wgmma Instructions for H800 Tensor Cores. Latency (LAT) is quantified in clock cycles, while throughput is expressed in TFLOPS or TOPS. The peak throughputs for FP16, TF32, FP8, and INT8 are 756.5, 373, 1513, and 1513, correspondingly. “Zero” or “Rand” signifies that all matrices are initialized with either zero or randomly generated values. “SS” implies that both matrix A and B are stored in shared memory, while “RS” signifies that matrix A is stored in the register file, whereas B is stored in shared memory. “Rand” has the same latency as “Zero”.
@@
In the context of sparse wgmma instructions, the latencies for “RS” and “SS” modes are 128.0 and 144.0, respectively. Additionally, we observe that in the ”SS” mode (Rand), sparse wgmma achieves approximately 1.8x performance improvement compared to dense wgmma. In the ”RS” mode (Rand), the performance improvement exceeds 1.9x. However, the performance of sparse wgmma in the ”SS” mode is lower than in the ”RS” mode, which is notably different from the behavior of dense wgmma. We find that in sparse wgmma, the “SS” mode retrieves data from the shared memory of size $m\times k$ and performs a 2:4 sparse pruning based on metadata during the execution of the sparse wgmma instruction. In contrast, the “RS” mode directly accesses data from the pruned register file of size $m\times k/2$. The high shared memory access demand (twice as much) may lead to latency that cannot be effectively concealed by Tensor Core computation, resulting in sparse wgmma instructions in “SS” modes failing to achieve the expected peak performance.
wgmma results with different $N$ values. We conduct tests using the example of wgmma.m64nNk16.f32.f16.f16, varying the value of $N$, and the results are presented in Table 9. When $N$ is greater than or equal to 64, all wgmma instructions can achieve throughputs that closely approach peak performance. However, when $N$ is less than 64, the achieved throughput decreases, and the “SS” mode of instructions exhibits higher latency than the “RS” mode, while the achieved throughputs are lower than those of the “RS” mode. As $N$ decreases, the computational density of wgmma instructions gradually diminishes, making it challenging to conceal the latency associated with shared memory access, leading to the aforementioned phenomena. Therefore, when utilizing wgmma instructions, it is advisable to opt for larger values of NN (>=64>=64) whenever possible to attain superior performance.
Table 9. Different wgmma instructions with different NN values on H800 tensor cores. The definitions of LAT and throughput can be found in the caption of Table 8. “Rand” has the same latency as “Zero”.
N Dense Sparse
LAT/Tput
($S$,Zero)
LAT/Tput
@@
Tput
($S$,Rand)
Tput
@@
LAT/Tput
($S$,Zero)
LAT/Tput
@@
Tput
($S$,Rand)
Tput
@@
Energy efficiency. Although Tensor Cores have impressive performance, their energy efficiency should also be considered. The training of GPT-3 consumes approximately 1,287 megawatt-hours of electricity in a single training run(Maslej et al., 2023). Since tensor core instructions constitute the primary computational workload in AI training, analyzing their energy consumption is essential for developing green AI solutions(Tang et al., 2019; Wang et al., 2020; Nabavinejad et al., 2022). Therefore, we conduct comprehensive energy efficiency analysis to understand the power characteristics of different tensor core instructions.
@@
In our benchmarks, we define energy efficiency as performance per watt, measured in TFLOPS(TOPS)/watt. Both mma and wgmma are considered in this work. Our methodology involves repeatedly executing either mma or wgmma instructions across all SMs, with a minimum of two billion repetitions. We utilize the tools provided in (Wang and Chu, 2020) to monitor both frequency and performance. The energy efficiency of each instruction is determined by calculating the ratio of performance to power consumption once steady-state operation is achieved during these repetitive executions.
@@
For mma instructions, we test the largest operational shape in Table 7. The energy efficiency of mma instructions is shown in Table 10. In terms of dense instructions, the average energy efficiency of H800 is 1.60 times and 1.69 times that of A100 and RTX4090 respectively. In terms of sparse instructions, the average energy efficiency of H800 is 1.33 times and 1.39 times that of A100 and RTX4090 respectively. We find that the H800 has significantly higher energy efficiency.
@@
Table 10. Power consumption and energy efficiency of maximum shape under mma instructions. Energy is measured in Watts. Efficiency is measured in TFLOPS(TOPS)/Watt.
A/B C/D Type A100 H800 RTX4090
Energy Efficiency Energy Efficiency Energy Efficiency
FP16 FP16 Dense 173.4 1.79 188.6 2.62 189.1 1.89
Sparse 198.8 3.13 187.2 3.86 214.0 3.33
FP16 FP32 Dense 188.5 1.61 196.7 2.49 154.1 1.16
Sparse 216.1 2.79 194.9 3.70 165.9 2.15
TF32 FP32 Dense 214.7 0.71 254.9 0.97 174.3 0.51
Sparse 235.7 1.28 232.5 1.56 187.9 0.95
INT8 INT32 Dense 178.4 3.41 165.3 5.92 201.4 3.53
Sparse 193.9 6.24 163.3 8.79 219.8 6.47
@@
For wgmma instructions, we focus on both the frequency and efficiency when these instructions are executed continuously. The instruction formats are consistent with those described in Table 8. The results are presented in Table 11. It is observed that, during continuous execution of wgmma instructions, the core frequency drops below 1620 MHz specified in the whitepaper, as it reaches the 350-watt power limit. Furthermore, the initialization of inputs significantly impacts both power consumption and frequency. When inputs are set to zero, power consumption remains below 200 watts. However, with random inputs, the power consumption rapidly hits the 350-watt threshold. Consequently, the performance results in Table 11 are inferior to those in Table 8. Notably, sparse instructions cause a substantial reduction in frequency, explaining why their performance is not twice that of dense instructions. Regarding energy efficiency, wgmma dense and sparse instructions achieve average 0.67 times and 0.78 times the efficiency of mma dense and sparse instructions, respectively. Compared to different architectures, the energy efficiency of Hopper wgmma is on average 1.05 times and 1.10 times that of Ampere mma and Ada mma instructions, respectively.